(A5). A random variable X has Poisson distribution with mean equal to 0.4. What is the probability that the random variable is greater than zero?
1
Expert's answer
2020-06-02T18:58:20-0400

"\lambda =0.4\\
P(X>0)=1-P(X=0)\\
=1-\frac{e^{-0.4}(0.4)^{0}}{0!}=1-0.0.67=0.33"